# Dissipative qutrit-mediated stable charging

### v1

@@ III Effective master equation
$\mathcal{E}(t)\equiv{\rm Tr}[H_{B}(\rho_{B}(t))]-{\rm Tr}[H_{B}\sigma(t)],$ (18)
@@
Here $\sigma(t)$ is the passive state that is diagonal in the battery eigenstates basis with non-increasing eigenvalues of $\rho_{B}(t)$:
$\sigma(t)=\sum_{n=0}^{N}\lambda_{n}(t)|n\rangle\langle n|,\quad\lambda_{n+1}(t)\leq\lambda_{n}(t)$ (19)
@@
Both energy and ergotropy saturate when the battery state approaches the maximal-ergotropy state $\rho_{B}\rightarrow|N\rangle\langle N|$. By virtue of the last term in Eq. (8) interpreted by Eq. (10b), our charging scheme can hold on that state without decoupling the charger from the battery after the charging process is completed.
@@
In Fig. 3, we compare the dynamics of the battery energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ under the full master equation (4) and the effective master equation (8). The battery is initially prepared at its ground state $\rho_{B}(0)=|0\rangle\langle 0|$ with vanishing energy and ergotropy. It is found that for either energy in Fig. 3(a) or ergotropy in Fig. 3(b), the charging efficiency becomes higher under a stronger driving intensity $\Omega$. On the other hand, the deviation between the effective dynamics and the full dynamics increases when $\Omega$ becomes larger and larger. It is due to the breakdown of the weak driving assumption.
@@
A practical charging scheme can transform a high-energy yet zero-ergotropy state, e.g., the thermal state, to be a high-energy and high-ergotropy state. In Fig. 4(a) and 4(b), we demonstrate the charging performance for the initial thermal states with various inverse temperatures $\beta$ by the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$, respectively. $\beta=\infty$ means the ground state. It can be found that both energy and ergotropy still monotonically increase with the charging time and ultimately stabilise with their respective maximum values. Note high-ergotropy state is different from high-energy state in population distribution. Due to the unidirectional transition from the lower Fock states to the higher ones as indicated in Eq. (10b), an initially higher-temperature battery state is more convenient to be transformed to be a higher-ergotropy state than the lower-temperature state. The final increment of $\Delta E$ for an initially high-temperature state is significantly smaller than that for $\mathcal{E}(t)$. Moveover, we take account of a nonvanishing $\gamma_{he}$ in Fig. 4. In comparison to Fig. 3, one can find that the effect of the process shown in Fig. 2(b) is marginal in the charging efficiency, that is consistent with our analysis.
@@ IV Large spin battery
Figure 6: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under the optimization conditions of $g=g_{m=-J,{\rm opt}}$ (solid lines and dashed lines) and $g=g_{m=0,{\rm opt}}$ (dotted lines and dot-dashed line), as defined in Eq. (23). The quantum number of energy is set as $J=25$ and the decay rate is $\gamma_{he}=\gamma_{eg}$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
To implement our charging scheme in a more physically relevant scenario, we consider an $N=2J+1$ dimensional large-spin battery in this section. The battery Hamiltonian reads $H_{B}=E_{B}J_{z}=E_{B}\sum_{m=-J}^{J}m|J,m\rangle\langle J,m|$, where $J$ is an integer or half-integer quantum number of energy. The charger and the battery interact with each other through
$H_{I}=g(J_{+}|e\rangle\langle h|+J_{-}|h\rangle\langle e|),$ (20)
where the creation operator of battery is the collective angular momentum operator
@@
In contrast to the uniform model with $A_{n}=1$, the effective transition rates in the large spin model vary across different subspaces, the same is true for the optimization condition:
$g_{m,{\rm opt}}^{2}[J(J+1)-m(m+1)]=|\tilde{\Delta}\tilde{\delta}|.$ (23)
@@
Figure 7: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with different inverse temperature $\beta$. The other parameters are the same as in Fig. 6.
@@
In Fig. 5, we present the dependence of effective transition rate $\gamma_{m,{\rm eff}}$ on the coupling strength between the qutrit charger and the large-spin battery $g$ and the quantum number of angular momentum $m$ to confirm the optimization condition in Eq. (23), as indicated by the black dashed line. It is noted that by setting $m=0$ in Eq. (23), the relevant $\gamma_{m,{\rm eff}}$ closely approximates the optimal conditions across a substantial number of subspaces as shown by the red dash-dotted line. In practice, it is a $m$-independent setting for enhancing the charging efficiency.
In Fig. 6(a) and 6(b), we compare the charging efficiency under the optimized condition of $g=g_{m=-J,{\rm opt}}$ with that under $g=g_{m=0,{\rm opt}}$ when the battery is initialised as the ground state $\rho_{B}(0)=|J,-J\rangle\langle J,-J|$. It is found that while the optimized condition $g=g_{-J,{\rm opt}}$ exhibits an advantage at the beginning of charging due to the initial population distribution, the charging rate under $g=g_{0,{\rm opt}}$ dramatically increases as the charging proceeds. The latter quickly surpasses the former. This behavior is consistent with our effective master equation. Nevertheless, the effective dynamics under $g=g_{0,{\rm opt}}$ exhibits a moderate deviation from the full one, and it perfectly matches with the full dynamics under $g=g_{-J,{\rm opt}}$.
@@
Now we check the case that the large-spin battery is initially prepared at the thermal states with various inverse temperature $\beta$ by the dynamics of stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ in Fig. 7(a) and 7(b), respectively. The optimization condition is chosen as $g=g_{0,{\rm opt}}$. It is found that the ergotropy can be maintained about $\mathcal{E}_{\rm steady}>0.995\mathcal{E}_{\rm max}$, irrespective to the temperature $\beta$ of the initial state, where $\mathcal{E}_{\rm max}=2JE_{B}$. Moreover, it is interesting to find that a higher initial temperature of the battery results in a higher growth rate of the ergotropy.
@@ V Harmonic oscillator battery
$g_{n,{\rm opt}}^{2}(n+1)=|\tilde{\Delta}\tilde{\delta}|.$ (25)
Figure 9: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with an invariant coupling strength $g=g_{0,{\rm opt}}$ in Eq. (25) during the charging process (black solid lines and blue dashed lines) and with a coupling strength quenched at $\gamma_{eg}\tau$ (green and yellow lines) according to Eq. (26). The decay rate $\gamma_{he}=\gamma_{eg}$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
The effective transition rate in the parameter space of charger-battery coupling strength $g$ and $n$ is plotted in Fig. 8. The optimization condition is indicated by the black dashed line. Distinct from the large-spin battery, it is hardly to find a subspace-independent coupling strength $g_{\rm opt}$ to optimize the effective transition rates for the harmonic-oscillator battery, since $g_{n,{\rm opt}}$ decreases with $n$ in a monotonic way. Nevertheless, the charging efficiency can be enhanced by adjusting the coupling strength during the charging process. By our scheme, one can initially optimize the effective transition rate from $|0\rangle$ to $|1\rangle$ by choosing $g=g_{0,{\rm opt}}$ in Eq. (25) due to the fact that the ground state is always the most populated state for a thermal state. Then after an interval of $\tau$, one can modify the coupling strength to be
@@
We compare the charging performance with a uniform $g$ during the whole charging evolution and that with a quenched $g$ by the dynamics of $\Delta E(t)$ and $\mathcal{E}(t)$ in Figs. 9(a) and 9(b), respectively. The yellow and green lines indicate the quenched moment at $\gamma_{eg}\tau=500$ and $\gamma_{eg}\tau=1000$, respectively, in which the charger-battery coupling strength is only modified at the specified moment and calculated according to Eq. (26). The battery is initially prepared in the ground state $\rho_{B}(0)=|0\rangle\langle 0|$. It is clearly that the strategy of a uniform $g$ is the lowest in charging efficiency and both energy and ergotropy have an abrupt and significant enhancement when $g$ can be re-optimized. The numerical simulation indicates that the full dynamics is perfectly matched with the effective dynamics and the charging efficiency is slightly overestimated after the adjustment of the coupling strength $g$.
Figure 10: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various $\beta$. The coupling strength is only modified at $\gamma_{eg}\tau=500$ according to Eq. (26). The other parameters are the same as in Fig. 9.
The dynamics of the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ for charging the HO battery initially in the thermal states with a finite temperature is demonstrated in Figs. 10(a) and 10(b), respectively. The results are similar to the large-spin battery presented in Fig. 7. A higher initial temperature or energy results in a higher growth rate of the ergotropy. In addition, the steady value of the ergotropy is insensitive to the initial inverse temperature $\beta$. In this case, we have $\mathcal{E}_{\rm steady}>0.994\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=NE_{B}$.
@@ VI Conclusion
In summary, we have proposed a stable charging scheme assisted by a dissipative $\Delta$-type qutrit, which relies on the nonreciprocal interaction between the external power source and the battery and the ensued unidirectional energy flow. Three types of finite dimensional quantum batteries, including the uniform coupling strength model, the large-spin battery, and the truncated harmonic oscillator battery, have confirmed that the battery energy and ergotropy can be maintained at their respective maximum values and the battery need not to be decoupled from the charger after the charging process is completed. By deriving the effective master equation of the battery, we can obtain the optimization condition of the charging efficiency. For the large spin battery, the charging efficiency can be significantly enhanced by choosing the charger-battery coupling strength according to the optimization condition in the subspace with $m=0$. For the truncated harmonic oscillator battery, the enhancement can be achieved by adjusting the coupling strength according to the instantaneous state of the battery during charging. In application, our work provides an efficient initial-state-independent charging scheme and avoids the discharging induced by the time-reversal symmetry. In theory, it extends the high-order Fermi’s Golden rule to the regime of non-Hermitian Hamiltonian.
Appendix A The effective master equation
This appendix contributes to deriving the effective Hamiltonian and Lindblad operators for the quantum battery $\rho_{B}(t)$ using a technique extending the James’ method James and Jerke (2007); Gamel and James (2010); Shao et al. (2017) from Hamiltonian to Lindbladian.

### v2

@@ III Effective master equation
$\mathcal{E}(t)\equiv{\rm Tr}[H_{B}\rho_{B}(t)]-{\rm Tr}[H_{B}\sigma(t)].$ (19)
@@
Here $\sigma(t)$ is the passive state that is diagonal in the battery eigenbases with nonincreasing eigenvalues of $\rho_{B}(t)$:
$\sigma(t)=\sum_{n=0}^{N}\lambda_{n}(t)|n\rangle\langle n|,\quad\lambda_{n+1}(t)\leq\lambda_{n}(t).$ (20)
@@
Both energy and ergotropy saturate when the battery approaches the maximal-ergotropy state $\rho_{B}\rightarrow|N\rangle\langle N|$. By virtue of the last term in Eq. (8) or Eq. (10b), our charging scheme can hold that state without decoupling the charger from the battery after the charging is completed. The first and second terms in Eq. (8) do not contribute to the battery energy variation since both the effective Hamiltonian $H_{\rm eff}^{B}$ in Eq. (9) and the effective Lindblad operator $L^{B}_{hg,{\rm eff}}$ in Eq. (10a) are diagonal in the energy eigenstates.
Figure 3: Comparison between the dynamics obtained by the full master equation (4) and the effective master equation (8) of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with various $\Omega$. (c) Histogram of the population distribution of the battery eigenlevels $\langle n|\rho_{B}(t)|n\rangle$ at various moments when a certain amount of ergotropy $\mathcal{E}(t)$ has been accumulated in the battery under $\Omega=0.05\gamma_{hg}$. The coupling strength $g$ satisfies the optimized condition (16). The battery size is set as $N=50$. $\gamma_{he}=0$, $\gamma_{hg}=10\gamma_{eg}=0.1E_{B}$, and $\Delta=10\delta=0.1E_{B}$.
@@
In Figs. 3(a) and (b), we compare the dynamics of the battery energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ under the full master equation (4) and the effective master equation (8). The battery is initially prepared at its ground state $\rho_{B}(0)=|0\rangle\langle 0|$ with zero energy and zero ergotropy. It is found that for either energy in Fig. 3(a) or ergotropy in Fig. 3(b), the charging efficiency becomes higher under a stronger driving intensity $\Omega$. In the mean time, the deviation between the effective dynamics and the full dynamics is also enlarged with increasing $\Omega$, which is due to the breakdown of the weak-driving assumption. An interesting observation is a sudden slope change during the ergotropy dynamics, roughly around $\mathcal{E}(t)\sim 30E_{B}$ as shown in Fig. 3(b). In contrast, the energy increases smoothly during the charging.
As a unitary-extractable energy, the ergotropy is the gap between the energy of the battery state $\rho_{B}(t)$ and that of the passive state $\sigma(t)$. It is then more sensitive than energy to the variation of population distribution. Our numerical simulation shows that once the ergotropy exceeds $\mathcal{E}(t)\approx 30.8E_{B}$, the top energy level becomes the most populated one, indicating a dominant inversion in the population distribution. As shown in Fig. 3(c), at the moments when the battery is charged with $\mathcal{E}(t)=25E_{B}$, $30E_{B}$, and $35E_{B}$, the top-level populations are about $0.005$, $0.05$, and $0.15$, respectively. The dramatic raise of the top-level population renders a dramatic change in the eigenvalues of the battery density operator $\rho_{B}(t)$ as well as the population histogram. Due to the definition in Eq. (20), the passive-state energy will stop rising and turn to descend as the top eigenvalue of $\rho_{B}(t)$ has a dramatic increment. The ergotropy therefore experiences a subtle yet sudden change in slope around $\mathcal{E}(t)\sim 30E_{B}$. On the other hand, the stored energy continues to rise smoothly under charging as it is just the weighted sum of population distribution.
Figure 4: Charging performance with decay from $|h\rangle$ to $|e\rangle$, $\gamma_{he}=\gamma_{eg}$, under various inverse temperature $\beta$ ($k_{B}\equiv 1$). (a) Stored Energy $\Delta E(t)$. (b) Ergotropy $\mathcal{E}(t)$. $\Omega=0.05\gamma_{hg}$ and the other parameters are the same as in Fig. 3.
@@
A practical charging scheme can transform a high-energy yet zero-ergotropy state, e.g., the thermal state, to be a high-energy and high-ergotropy state. In Figs. 4(a) and (b), we demonstrate the charging performance for the initial thermal states with various inverse temperatures $\beta$. $\beta=\infty$ means the ground state. Both energy and ergotropy still monotonically increase and eventually hold on the maximum value. Note high-ergotropy states are different from high-temperature states in population distribution. The final increment of $\Delta E$ for an initially high-temperature state is significantly smaller than that of $\mathcal{E}(t)$. Due to the unidirectional transition from the lower to the higher states as described by Eq. (10b), an initially higher-temperature battery state is more convenient to be transformed to a higher-ergotropy state than the lower-temperature state. The two-stage behavior of the ergotropy in slope becomes gradually subtle as the battery starts from a higher temperature state [see the yellow and green lines of Fig. 4(b)]. Due to the fact that from the beginning, the top level is significantly populated and then the change in the eigenvalues of $\rho_{B}(t)$ is not as dramatic as that from a zero-temperature state [see the blue lines of Fig. 4(b)]. In addition, we take account of a finite $\gamma_{he}$ in Fig. 4. In comparison to Fig. 3, one can find that the discharging depicted in Fig. 2(b) has a marginal effect on the charging efficiency, that is consistent with our analysis.
@@ IV Large spin battery
To implement our charging scheme in a more physically relevant scenario, we consider a $(2J+1)$-dimensional large-spin battery in this section. The battery Hamiltonian reads $H_{B}=E_{B}J_{z}=E_{B}\sum_{m=-J}^{J}m|J,m\rangle\langle J,m|$, where $J$ is an integer or half-integer quantum number of energy. The charger and the battery interact with each other through $H_{I}=g(J_{+}|e\rangle\langle h|+J_{-}|h\rangle\langle e|)$, where the creation operator of battery is the collective angular momentum operator
@@
In contrast to the uniform model with $A_{n}=1$, the effective transition rates in the large spin model vary across different subspaces. The same is true for the optimization condition:
[…]
@@
Figure 6: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under the optimization conditions of $g=g_{m=-J,{\rm opt}}$ (solid lines and dashed lines) and $g=g_{m=0,{\rm opt}}$ (dotted lines and dot-dashed lines), as defined in Eq. (23). $J=25$, $\gamma_{he}=\gamma_{eg}$, and $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
@@
In Fig. 5, we present the dependence of effective transition rate $\gamma_{m,{\rm eff}}$ on the coupling strength $g$ between the qutrit charger and the large-spin battery and the quantum number of angular momentum $m$. The optimization condition in Eq. (23) is indicated by the black dashed line. It is found that by setting $m=0$ in Eq. (23), the relevant $\gamma_{m,{\rm eff}}$ closely mimics the optimal conditions across a substantial number of subspaces as shown by the red dash-dotted line. In practice, it is an $m$-independent setting for enhancing the charging efficiency.
In Figs. 6(a) and 6(b), we compare the charging efficiency under the optimized condition of $g=g_{m=-J,{\rm opt}}$ with that under $g=g_{m=0,{\rm opt}}$ when the battery is initialized as the ground state $\rho_{B}(0)=|J,-J\rangle\langle J,-J|$. Under the optimized condition $g=g_{0,{\rm opt}}$, the ergotropy remains almost vanishing during the beginning stage of the charging, while the energy grows with a small rate. It is because the effective transition rate from $|J,-J\rangle$ to $|J,-J+1\rangle$ is much smaller than the optimized one from $|J,0\rangle$ to $|J,1\rangle$, i.e., the battery is almost in a passive state $\rho_{B}(t)\approx\sigma(t)$ on the beginning stage. As the charging proceeds, the ergotropy becomes nonvanishing when the population inversion around $|J,-J\rangle$ is realized. The optimized condition $g=g_{-J,{\rm opt}}$ exhibits an advantage in charging rate on the beginning stage over $g=g_{0,{\rm opt}}$ due to the initial population distribution, i.e., the population on $|J,-J\rangle$ can be efficiently moved to the high level $|J,-J+1\rangle$. After a sufficient amount of energy accumulated in the battery, the condition $g=g_{0,{\rm opt}}$ quickly exceeds $g=g_{-J,{\rm opt}}$ since a dominant number of subspaces are efficient in charging (see Fig. 5). Nevertheless, the effective dynamics under $g=g_{0,{\rm opt}}$ exhibits a moderate deviation from the full one and it matches perfectly with the full dynamics under $g=g_{-J,{\rm opt}}$. It is due to the tradeoff between the charging efficiency and the validity of the effective master equation under weak driving. A higher charging efficiency means the excited levels $|h\rangle$ and $|e\rangle$ of the charger are more populated as indicated by Fig. 2(a), that disrupts the approximation $\rho(t)\approx P_{g}\rho(t)P_{g}$. On the contrary, only the transition $|J,-J\rangle\rightarrow|J,-J+1\rangle$ is efficient under $g=g_{-J,{\rm opt}}$. The slow charging then supports the agreement between the effective and full master equations.
Figure 7: Full and effective dynamics of (a) energy stored in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various inverse temperatures $\beta$. The other parameters are the same as in Fig. 6.
@@
Now we check the large-spin battery prepared as finite-temperature states. The dynamics of stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ are plotted in Figs. 7(a) and (b), respectively. The optimization condition for charging is chosen as $g=g_{0,{\rm opt}}$. The steady ergotropy is held about $\mathcal{E}_{\rm steady}>0.995\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=2JE_{B}$, irrespective of $\beta$. Moreover, it is interesting to find again that a higher initial temperature of the battery results in a higher growth rate of the ergotropy.
@@ V Harmonic oscillator battery
[…]
The effective transition rate in the space of the coupling strength $g$ and the subspace index $n$ is plotted in Fig. 8. The optimization condition is indicated by the black dashed line. Distinct from the large-spin battery, it is hardly to find a nearly subspace-independent coupling strength $g_{\rm opt}$ to optimize the effective transition rate for the HO battery, since $g_{n,{\rm opt}}$ decreases monotonically with $n$. Nevertheless, the charging efficiency can be enhanced by quenching the coupling strength during the charging process. By our scheme, one can initially choose $g=g_{0,{\rm opt}}$ since the ground level is always the most occupied one for a thermal state. Then after a duration $\tau$, the coupling strength can be quenched as
@@
Figure 9: (a) Saturation time $\gamma_{eg}T$ of ergotropy versus the quench time $\gamma_{eg}\tau$. (b) Effective dynamics of the ergotropy $\mathcal{E}(t)$ under various settings of quenching. The coupling strength is initially $g=g_{0,{\rm opt}}$ in Eq. (25) and then quenched according to Eq. (26). The decay rate $\gamma_{he}=\gamma_{eg}$. $\Omega=0.1\gamma_{hg}$. The other parameters are the same as in Fig. 3.
In Fig. 9(a), we demonstrate the saturation time $\gamma_{eg}T$ for charging the truncated HO battery to the top energy level as a function of the quench time $\gamma_{eg}\tau$. It is found that $\gamma_{eg}T$ can have a comparably small and steady value in the range of $500\leq\gamma_{eg}\tau\leq 1000$. In addition, the saturation time of charging can be further reduced if one can apply multiple quenches within that range of quench time. As shown in Fig. 9(b), the quench effect is quickly saturated.
Figure 10: Full and effective dynamics of (a) battery energy $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ with a fixed coupling strength $g=g_{0,{\rm opt}}$ during the charging process (black solid lines and blue dashed lines) and with a coupling strength quenched different moments (green and yellow lines). The other parameters are the same as in Fig. 9.
Figure 11: Full and effective dynamics of (a) stored energy in the battery $\Delta E(t)$ and (b) ergotropy $\mathcal{E}(t)$ as functions of $\gamma_{eg}t$ under initial thermal states with various $\beta$. The coupling strength is only modified at $\gamma_{eg}\tau=500$ according to Eq. (26). The other parameters are the same as in Fig. 10.
We present the charging performance with a fixed $g$ and that with a quenched $g$ by the dynamics of $\Delta E(t)$ in Fig. 10(a) and $\mathcal{E}(t)$ in Fig. 10(b). The yellow and green lines indicate the quenched moment at $\gamma_{eg}\tau=500$ and $\gamma_{eg}\tau=1000$, respectively. The battery is initially prepared in the ground state $\rho_{B}(0)=|0\rangle\langle 0|$. Apparently, the dynamics of both energy and ergotropy have an abrupt and dramatic acceleration when $g$ can be re-optimized. The numerical simulation indicates that the full dynamics perfectly matches with the effective dynamics and the charging efficiency is slightly overestimated after the adjustment of the coupling strength $g$.
The dynamics of the stored energy $\Delta E(t)$ and ergotropy $\mathcal{E}(t)$ for charging the HO battery initially in the finite-temperature states is demonstrated in Figs. 11(a) and (b), respectively. The results are similar to the large-spin battery presented in Fig. 7. A higher initial temperature or energy results in a higher growth rate of the ergotropy. Again, the steady value of the ergotropy is insensitive to the initial inverse temperature $\beta$. In this case, we have $\mathcal{E}_{\rm steady}>0.994\mathcal{E}_{\rm max}$ with $\mathcal{E}_{\rm max}=NE_{B}$.
@@ VI Conclusion
In summary, we proposed a stable charging scheme assisted by a dissipative qutrit yielding an unidirectional energy flow, which is based on the effective nonreciprocal interaction between the external power source and the battery. Three prototypes of finite dimensional quantum batteries, including the uniform coupling-strength model, the large-spin battery, and the truncated harmonic oscillator battery, have confirmed that the battery energy and ergotropy can hold on the maximum values and the battery need not to be decoupled from the charger after the charging process is completed. By deriving the effective master equation for the battery, that could be alternatively obtained by the generalized Fermi’s golden rule adapted to non-Hermitian Hamiltonian and spontaneous decay, one can extract the optimized condition of the charging rate. For the large-spin battery, the charging efficiency can be significantly enhanced by choosing the charger-battery coupling strength according to the optimization condition in the subspace indexed with $m=0$. For the truncated harmonic oscillator battery, the enhancement can be achieved by quenching the coupling strength according to the instantaneous state of the battery. In application, our work provides an efficient initial-state-independent charging scheme and avoids the discharging induced by the time-reversal symmetry.
Acknowledgments
We acknowledge grant support from the Science and Technology Program of Zhejiang Province (No. 2025C01028).
Appendix A Effective master equation
This appendix contributes to deriving the effective Hamiltonian and Lindblad operators for the quantum battery $\rho_{B}(t)$ using a technique extending the James’ method James and Jerke (2007); Gamel and James (2010); Shao et al. (2017) from Hamiltonian to Lindbladian. The fundamental idea is still based on the large detuning condition and the relevant rotating-wave approximation.